in multiplying polynomials, what would be the answer to 2a+1 times a+6?
Answer by stanbon(75887)

in multiplying polynomials, what would be the answer to 2a+1 times a+6?
---
Use FOIL:
(2a+1)(a+6)
-------------
F: 2a*a = 2a^2
O: 2a*6 = 12a
I: 1*a = a
L: 1*6 = 6
-----
Combine: 2a^2+13a+6
====================
Cheers,
Stan H.
====================